# Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values

### v2

@@ Appendix H Extended Evaluation Results
Table 9: Full GSM8K Accuracies (Flexible).
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0.0106141 0.0121304 0.0159212 0.00985595 0.0159212 0.0106141 0.0144049 0.0204701
k1v2 0.00833965 0.0144049 0.015163 0.00985595 0.0197119 0.0128886 0.0174375 0.0181956
k1v4 0.0090978 0.0197119 0.025019 0.0144049 0.0204701 0.0075815 0.00985595 0.0106141
k1v6 0.00985595 0.0106141 0.0272934 0.0174375 0.0197119 0.0075815 0.0045489 0.0159212
k1v8 0.00833965 0.015163 0.0280516 0.0090978 0.015163 0.0136467 0.0113723 0.0106141
k2v1 0.0030326 0.0181956 0.015163 0.00833965 0.0500379 0.0128886 0.0204701 0.0144049
k2v2 0.0030326 0.0288097 0.278241 0.0265353 0.38514 0.128886 0.327521 0.38514
k2v4 0.0030326 0.0197119 0.351024 0.0432146 0.45489 0.170584 0.412434 0.43442
k2v6 0.00379075 0.0257771 0.3442 0.040182 0.445792 0.169067 0.396513 0.419257
k2v8 0.00379075 0.0280516 0.354814 0.0348749 0.460197 0.166793 0.394238 0.448825
k4v1 0.0227445 0.0727824 0.281274 0.576952 0.517058 0.77862 0.70887 0.721001
k4v2 0.040182 0.294162 0.639121 0.808188 0.756634 0.87718 0.804397 0.648976
k4v4 0.040182 0.332828 0.652767 0.830174 0.777862 0.87718 0.823351 0.629265
k4v6 0.055345 0.33207 0.66793 0.831691 0.762699 0.88097 0.828658 0.619409
k4v8 0.0416983 0.341168 0.658074 0.826384 0.75815 0.877938 0.835481 0.631539
k6v1 0.0932525 0.0962851 0.307809 0.632297 0.521607 0.795299 0.742987 0.701289
k6v2 0.367703 0.3116 0.644428 0.830174 0.761941 0.880212 0.824867 0.59818
k6v4 0.404852 0.333586 0.653525 0.836998 0.770281 0.877938 0.827142 0.610311
k6v6 0.420015 0.336619 0.662623 0.843821 0.771797 0.884003 0.8279 0.595906
k6v8 0.41395 0.341926 0.664898 0.840788 0.771039 0.886277 0.826384 0.608036
k8v1 0.0879454 0.0955269 0.304018 0.636846 0.529947 0.797574 0.742987 0.708112
k8v2 0.389689 0.313116 0.64746 0.825625 0.74981 0.882487 0.822593 0.605762
k8v4 0.416224 0.340409 0.660349 0.832449 0.770281 0.874905 0.831691 0.620925
k8v6 0.419257 0.341926 0.667172 0.843063 0.767248 0.880212 0.839272 0.608036
k8v8 0.413192 0.333586 0.666414 0.843063 0.758908 0.880212 0.839272 0.603487
bf16 0.412434 0.328279 0.663381 0.845337 0.759666 0.87718 0.835481 0.603487
Table 10: Full GSM8K Accuracies (Strict).
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0 0 0 0 0 0 0 0
k1v2 0 0 0 0 0 0 0 0
k1v4 0 0 0 0 0 0 0.00075815 0
k1v6 0 0 0.0015163 0 0 0.00075815 0 0
k1v8 0 0 0 0 0 0 0.00075815 0
k2v1 0 0.00075815 0.00530705 0 0.0280516 0 0.00075815 0.0015163
k2v2 0 0.0113723 0.275967 0.00379075 0.379833 0.0667172 0.260045 0.234268
k2v4 0 0.00985595 0.350265 0.0090978 0.443518 0.103867 0.343442 0.334344
k2v6 0 0.0121304 0.3442 0.00985595 0.438969 0.115239 0.348749 0.329795
k2v8 0 0.0136467 0.353298 0.00530705 0.448825 0.109174 0.35254 0.347233
k4v1 0.0060652 0.0614102 0.292646 0.64746 0.500379 0.785444 0.824867 0.701289
k4v2 0.0424564 0.297195 0.627748 0.824867 0.750569 0.875663 0.882487 0.726308
k4v4 0.0439727 0.331312 0.642911 0.843821 0.758908 0.875663 0.877938 0.733889
k4v6 0.055345 0.331312 0.656558 0.839272 0.747536 0.880212 0.887036 0.733131
k4v8 0.0485216 0.341926 0.648976 0.834723 0.745262 0.87718 0.887036 0.736922
k6v1 0.0833965 0.0826384 0.326005 0.698256 0.505686 0.802881 0.830174 0.702047
k6v2 0.369219 0.307809 0.639121 0.839272 0.74981 0.879454 0.884003 0.70887
k6v4 0.404852 0.333586 0.643669 0.844579 0.755876 0.876422 0.884761 0.718726
k6v6 0.421531 0.335102 0.6558 0.852919 0.751327 0.881729 0.884003 0.71721
k6v8 0.41395 0.341168 0.657316 0.84837 0.752843 0.884761 0.884003 0.721001
k8v1 0.0849128 0.0841547 0.32373 0.704321 0.512509 0.805155 0.829416 0.695224
k8v2 0.388173 0.313874 0.639879 0.829416 0.739196 0.881729 0.886277 0.71721
k8v4 0.416983 0.340409 0.652009 0.845337 0.755876 0.874905 0.88097 0.721001
k8v6 0.419257 0.339651 0.661107 0.84837 0.749052 0.878696 0.886277 0.722517
k8v8 0.415466 0.33207 0.659591 0.850644 0.743745 0.876422 0.886277 0.723275
bf16 0.413192 0.327521 0.657316 0.852161 0.741471 0.874147 0.890826 0.717968

### v3

@@ C.5 Downstream Accuracy Across Quantization Precisions
[…]
Qwen3 Llama-3.2 Llama-3.2 Qwen3 Llama-3.1 Qwen3 Qwen3
0.6B 1B 3B 4B 8B 8B 32B
$\mathbf{K}_{1}\mathbf{V}_{1}$ 0.011 0.012 0.016 0.010 0.016 0.011 0.020
$\mathbf{K}_{1}\mathbf{V}_{2}$ 0.008 0.014 0.015 0.010 0.020 0.013 0.018
$\mathbf{K}_{1}\mathbf{V}_{4}$ 0.009 0.020 0.025 0.014 0.020 0.008 0.011
$\mathbf{K}_{1}\mathbf{V}_{6}$ 0.010 0.011 0.027 0.017 0.020 0.008 0.016
$\mathbf{K}_{1}\mathbf{V}_{8}$ 0.008 0.015 0.028 0.009 0.015 0.014 0.011
$\mathbf{K}_{2}\mathbf{V}_{1}$ 0.003 0.018 0.015 0.008 0.050 0.013 0.014
$\mathbf{K}_{2}\mathbf{V}_{2}$ 0.003 0.029 0.278 0.027 0.385 0.129 0.385
$\mathbf{K}_{2}\mathbf{V}_{4}$ 0.003 0.020 0.351 0.043 0.455 0.171 0.434
$\mathbf{K}_{2}\mathbf{V}_{6}$ 0.004 0.026 0.344 0.040 0.446 0.169 0.419
$\mathbf{K}_{2}\mathbf{V}_{8}$ 0.004 0.028 0.355 0.035 0.460 0.167 0.449
$\mathbf{K}_{4}\mathbf{V}_{1}$ 0.023 0.073 0.281 0.577 0.517 0.779 0.721
$\mathbf{K}_{4}\mathbf{V}_{2}$ 0.040 0.294 0.639 0.808 0.757 0.877 0.649
$\mathbf{K}_{4}\mathbf{V}_{4}$ 0.040 0.333 0.653 0.830 0.778 0.877 0.629
$\mathbf{K}_{4}\mathbf{V}_{6}$ 0.055 0.332 0.668 0.832 0.763 0.881 0.619
$\mathbf{K}_{4}\mathbf{V}_{8}$ 0.042 0.341 0.658 0.826 0.758 0.878 0.632
$\mathbf{K}_{6}\mathbf{V}_{1}$ 0.093 0.096 0.308 0.632 0.522 0.795 0.701
$\mathbf{K}_{6}\mathbf{V}_{2}$ 0.368 0.312 0.644 0.830 0.762 0.880 0.598
$\mathbf{K}_{6}\mathbf{V}_{4}$ 0.405 0.334 0.654 0.837 0.770 0.878 0.610
$\mathbf{K}_{6}\mathbf{V}_{6}$ 0.420 0.337 0.663 0.844 0.772 0.884 0.596
$\mathbf{K}_{6}\mathbf{V}_{8}$ 0.414 0.342 0.665 0.841 0.771 0.886 0.608
$\mathbf{K}_{8}\mathbf{V}_{1}$ 0.088 0.096 0.304 0.637 0.530 0.798 0.708
$\mathbf{K}_{8}\mathbf{V}_{2}$ 0.390 0.313 0.647 0.826 0.750 0.882 0.606
$\mathbf{K}_{8}\mathbf{V}_{4}$ 0.416 0.340 0.660 0.832 0.770 0.875 0.621
$\mathbf{K}_{8}\mathbf{V}_{6}$ 0.419 0.342 0.667 0.843 0.767 0.880 0.608
$\mathbf{K}_{8}\mathbf{V}_{8}$ 0.413 0.334 0.666 0.843 0.759 0.880 0.603
BF16 0.412 0.328 0.663 0.845 0.760 0.877 0.603
[…]
Qwen3 Llama-3.2 Llama-3.2 Qwen3 Llama-3.1 Qwen3 Qwen3
0.6B 1B 3B 4B 8B 8B 32B
$\mathbf{K}_{1}\mathbf{V}_{1}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\mathbf{K}_{1}\mathbf{V}_{2}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\mathbf{K}_{1}\mathbf{V}_{4}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\mathbf{K}_{1}\mathbf{V}_{6}$ 0.000 0.000 0.002 0.000 0.000 0.001 0.000
$\mathbf{K}_{1}\mathbf{V}_{8}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\mathbf{K}_{2}\mathbf{V}_{1}$ 0.000 0.001 0.005 0.000 0.028 0.000 0.002
$\mathbf{K}_{2}\mathbf{V}_{2}$ 0.000 0.011 0.276 0.004 0.380 0.067 0.234
$\mathbf{K}_{2}\mathbf{V}_{4}$ 0.000 0.010 0.350 0.009 0.444 0.104 0.334
$\mathbf{K}_{2}\mathbf{V}_{6}$ 0.000 0.012 0.344 0.010 0.439 0.115 0.330
$\mathbf{K}_{2}\mathbf{V}_{8}$ 0.000 0.014 0.353 0.005 0.449 0.109 0.347
$\mathbf{K}_{4}\mathbf{V}_{1}$ 0.006 0.061 0.293 0.647 0.500 0.785 0.701
$\mathbf{K}_{4}\mathbf{V}_{2}$ 0.042 0.297 0.628 0.825 0.751 0.876 0.726
$\mathbf{K}_{4}\mathbf{V}_{4}$ 0.044 0.331 0.643 0.844 0.759 0.876 0.734
$\mathbf{K}_{4}\mathbf{V}_{6}$ 0.055 0.331 0.657 0.839 0.748 0.880 0.733
$\mathbf{K}_{4}\mathbf{V}_{8}$ 0.049 0.342 0.649 0.835 0.745 0.877 0.737
$\mathbf{K}_{6}\mathbf{V}_{1}$ 0.083 0.083 0.326 0.698 0.506 0.803 0.702
$\mathbf{K}_{6}\mathbf{V}_{2}$ 0.369 0.308 0.639 0.839 0.750 0.879 0.709
$\mathbf{K}_{6}\mathbf{V}_{4}$ 0.405 0.334 0.644 0.845 0.756 0.876 0.719
$\mathbf{K}_{6}\mathbf{V}_{6}$ 0.422 0.335 0.656 0.853 0.751 0.882 0.717
$\mathbf{K}_{6}\mathbf{V}_{8}$ 0.414 0.341 0.657 0.848 0.753 0.885 0.721
$\mathbf{K}_{8}\mathbf{V}_{1}$ 0.085 0.084 0.324 0.704 0.513 0.805 0.695
$\mathbf{K}_{8}\mathbf{V}_{2}$ 0.388 0.314 0.640 0.829 0.739 0.882 0.717
$\mathbf{K}_{8}\mathbf{V}_{4}$ 0.417 0.340 0.652 0.845 0.756 0.875 0.721
$\mathbf{K}_{8}\mathbf{V}_{6}$ 0.419 0.340 0.661 0.848 0.749 0.879 0.723
$\mathbf{K}_{8}\mathbf{V}_{8}$ 0.415 0.332 0.660 0.851 0.744 0.876 0.723
BF16 0.413 0.328 0.657 0.852 0.741 0.874 0.718
C.6 Rotation and Grouping Results
Tables 11-14 present the full set of downstream evaluation results for integrating rotation-based outlier redistribution with mixed-precision KV quantization across multiple models and tasks, including GSM8K (flexible and strict exact match), CoQA (F1), and CoQA (exact match). Each table reports accuracy for four key-value bit allocations ($\mathbf{K}_{2}\mathbf{V}_{2}$, $\mathbf{K}_{2}\mathbf{V}_{4}$, $\mathbf{K}_{4}\mathbf{V}_{2}$, $\mathbf{K}_{4}\mathbf{V}_{4}$) under four Rotations: none, key-only, value-only, and both key and value. The group size is fixed at 64 for both keys and values in these experiments to isolate the effect of rotation.
The results reveal several consistent trends across model scales and tasks. First, applying rotation to keys consistently improves performance at lower bit-widths ($\mathrm{K}_{2}\mathrm{V}_{2}$ and $\mathrm{K}_{2}\mathrm{V}_{4}$), significantly narrowing the gap to higher-precision baselines. In contrast, rotating values alone yields marginal or even negative effects, especially at low precisions. Notably, applying rotation to keys only under the $\mathrm{K}_{4}\mathrm{V}_{2}$ configuration achieves accuracy that closely matches the full $\mathrm{K}_{4}\mathrm{V}_{4}$ baseline, indicating that the dominant benefits of rotation stem from mitigating key outliers rather than value outliers. Applying rotation to both keys and values provides little additional gain beyond key-only rotation, reinforcing that keys are the primary target for rotation-based improvements.
Figure 11 examines the effect of group size configuration under a fixed $\mathrm{K}_{4}\mathrm{V}_{2}$ mixed-precision setting without rotation. Group size ($\mathrm{gs}$) determines the block granularity of quantization: smaller groups offer finer scaling at the cost of increased metadata and computation. The results reveal a clear asymmetry between keys and values. Reducing group size for keys consistently improves downstream accuracy across CoQA, GSM8K, EQ-Bench, and LongBench, with $\mathrm{gs}_{K}=32$ achieving the best overall performance. This reflects the higher sensitivity of key caches to quantization distortion. In contrast, increasing value group size to 64 or 128 has a negligible impact on accuracy while reducing overhead, consistent with their lower sensitivity.
Together, these findings provide practical guidance for combining geometry-driven bit allocation with rotation and grouping strategies: rotation should primarily target keys, while keys benefit from finer group granularity and higher precision; values, on the other hand, can use coarser grouping and lower precision with minimal accuracy loss.
Table 11: Rotation-based Outlier Redistribution Integration on GSM8K - Exact Match (Flexible). Downstream accuracy under different key-value bit allocations and rotation scopes. Group size is fixed to 64 for both keys and values.
Model Rotation $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
Llama 3.1-8B K+V 52.99 55.80 76.50 76.95
K only 52.92 57.32 75.44 76.50
V only 20.70 25.70 75.28 76.95
none 18.65 24.41 75.21 76.95
Llama 3.2-1B K+V 1.97 2.43 30.93 33.13
K only 2.05 3.03 28.43 33.13
V only 2.05 1.29 28.58 29.57
none 1.82 2.20 27.22 29.57
Llama 3.2-3B K+V 40.11 44.12 62.85 64.37
K only 38.06 45.26 61.18 63.91
V only 16.60 19.26 63.46 63.91
none 14.86 21.08 62.93 63.53
Qwen 0.6B K+V 0.76 0.61 37.83 36.47
K only 1.52 1.06 36.92 37.68
V only 1.52 1.67 1.36 2.12
none 0.83 1.90 1.36 2.05
Qwen 32B K+V 31.69 33.81 66.49 64.22
K only 30.10 33.51 63.99 65.28
V only 6.90 11.07 65.81 63.68
none 6.67 10.92 64.82 63.76
Qwen 4B K+V 0.91 0.76 83.78 85.37
K only 1.14 0.68 82.26 84.76
V only 0.83 1.14 82.56 83.70
none 1.14 0.83 81.80 83.40
Qwen 8B K+V 29.34 36.01 88.17 88.17
K only 29.57 35.56 87.64 87.95
V only 1.97 1.21 86.73 87.34
none 2.27 1.67 86.58 88.02
Table 12: Rotation-based Outlier Redistribution Integration on GSM8K - Exact Match (Strict). Downstream accuracy under different key-value bit allocations and rotation scopes. Group size is fixed to 64 for both keys and values.
Model Rotation $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
Llama 3.1-8B K+V 51.71 55.27 74.37 74.45
K only 51.93 56.48 73.24 73.92
V only 19.33 24.49 73.01 74.60
none 17.97 22.74 72.93 74.60
Llama 3.2-1B K+V 0.83 1.21 31.01 32.98
K only 0.68 1.29 29.04 33.06
V only 0.38 0.61 28.51 29.57
none 0.45 0.61 27.22 29.34
Llama 3.2-3B K+V 39.65 43.67 61.94 63.84
K only 36.47 44.96 60.27 63.46
V only 16.00 19.26 63.00 63.38
none 14.40 21.00 62.02 62.85
Qwen 0.6B K+V 0.00 0.00 38.14 36.92
K only 0.00 0.00 37.38 37.83
V only 0.00 0.00 0.15 0.53
none 0.00 0.00 0.30 0.08
Qwen 32B K+V 21.99 28.35 74.75 75.59
K only 21.46 26.38 74.75 75.28
V only 2.20 5.00 74.00 75.13
none 1.52 4.93 73.77 74.37
Qwen 4B K+V 0.23 0.08 84.53 85.60
K only 0.00 0.08 83.40 84.99
V only 0.00 0.00 82.49 83.70
none 0.00 0.08 81.80 83.78
Qwen 8B K+V 25.32 30.33 87.72 87.87
K only 25.17 32.15 87.49 87.34
V only 0.15 0.00 86.43 86.88
none 0.00 0.08 86.50 87.41
Table 13: Rotation-based Outlier Redistribution Integration on CoQA - F1 Score. Downstream accuracy under different key-value bit allocations and rotation scopes. Group size is fixed to 64 for both keys and values.
Model Rotation $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
Llama 3.1-8B K+V 77.70 77.83 78.88 78.45
K only 77.39 78.10 77.93 78.75
V only 75.46 76.71 79.34 78.76
none 74.72 76.57 78.10 79.18
Llama 3.2-1B K+V 50.66 53.55 70.44 70.47
K only 50.18 53.46 69.40 70.27
V only 39.50 44.93 69.58 70.52
none 38.82 44.33 69.53 70.31
Llama 3.2-3B K+V 78.45 78.78 79.24 79.13
K only 78.24 78.88 78.71 79.20
V only 73.25 74.48 78.86 79.27
none 74.10 74.55 79.14 79.37
Qwen 0.6B K+V 33.99 34.06 68.92 70.09
K only 33.88 33.70 69.62 70.08
V only 29.00 27.39 48.19 48.06
none 28.89 28.04 46.49 48.41
Qwen 32B K+V 64.85 67.39 82.46 82.59
K only 67.16 67.18 82.62 82.55
V only 77.82 78.55 82.35 82.50
none 76.80 78.35 82.60 82.77
Qwen 4B K+V 62.43 63.82 80.16 80.56
K only 62.09 64.35 80.46 80.63
V only 56.73 59.26 80.07 80.16
none 54.93 59.68 80.41 80.19
Qwen 8B K+V 79.33 79.65 82.32 81.98
K only 78.27 79.89 81.15 82.09
V only 67.51 70.41 82.23 82.01
none 66.86 69.83 81.49 81.99
Table 14: Rotation-based Outlier Redistribution Integration on CoQA - Exact Match. Downstream accuracy under different key-value bit allocations and rotation scopes. Group size is fixed to 64 for both keys and values.
Model Rotation $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
Llama 3.1-8B K+V 62.43 63.37 63.25 63.40
K only 62.12 63.32 63.03 63.65
V only 58.13 60.12 64.48 63.75
none 58.12 59.65 62.92 64.45
Llama 3.2-1B K+V 33.38 37.87 54.78 56.02
K only 33.72 37.30 54.22 55.32
V only 25.47 30.37 53.97 55.65
none 24.42 30.23 54.30 55.45
Llama 3.2-3B K+V 62.47 62.28 62.83 63.40
K only 62.12 62.72 62.93 63.53
V only 56.07 57.48 62.70 63.73
none 56.45 57.62 63.42 63.77
Qwen 0.6B K+V 19.67 19.57 56.07 56.73
K only 18.72 18.18 55.53 56.73
V only 13.97 14.27 28.18 28.17
none 14.25 13.80 28.22 28.82
Qwen 32B K+V 46.97 49.95 70.18 70.77
K only 50.28 49.42 70.48 70.87
V only 61.77 64.08 70.02 70.37
none 61.12 63.92 69.90 70.50
Qwen 4B K+V 41.28 42.45 65.68 66.28
K only 41.65 43.45 66.40 66.47
V only 39.22 40.03 65.43 66.20
none 37.70 40.22 65.52 66.07
Qwen 8B K+V 61.70 61.88 68.55 67.73
K only 60.07 62.40 66.62 68.23
V only 49.62 51.58 68.10 68.07
none 48.40 51.65 67.07 68.00
Figure 11: Effect of group size configurations on downstream accuracy. Each bar shows performance on CoQA, GSM8K, EQ-Bench, and LongBench for different key-value grouping combinations ($\mathrm{gs}_{K},\mathrm{gs}_{V}\in\{32,64,128\}$) under the $\mathbf{K}_{4}\mathbf{V}_{2}$ mixed-precision setting. Smaller group sizes for keys consistently improve accuracy across all tasks, while value group size has a smaller but non-negligible effect. Configurations with $\mathrm{gs}_{K}=32$ achieve the best overall performance, highlighting the benefit of finer quantization granularity for key caches, which are more sensitive to quantization distortion. In contrast, larger group sizes for values ($\mathrm{gs}_{V}=64$ or 128128) preserve performance while reducing overhead, aligning with their lower sensitivity.
Table 15: Performance by rotation strategy across tasks. Values report the mean performance across seven models (0.6B–32B parameters) and four quantization configurations ($\mathrm{K}_{2}\mathrm{V}_{2}$, $\mathrm{K}_{2}\mathrm{V}_{4}$, $\mathrm{K}_{4}\mathrm{V}_{2}$, $\mathrm{K}_{4}\mathrm{V}_{4}$). The $\pm$ values indicate the variability range across model scales and quantization settings, with larger ranges reflecting higher sensitivity to these factors.
Task Metric No Rotation Value-Only Key-Only Both
CoQa Exact Match $50.10\pm 17.65$ $1.82\pm 17.22 55$$34\pm 14.21 55.$$5.39\pm 14.10$
Eq Bench Parseable $6.18\pm 42.57 55$$5.01\pm 41.93$ $65.73\pm 39.70$ $64.81\pm 39.19$
GSM8K Exact Match $32.73\pm 33.42$ $32.96\pm 33.48$ $43.46\pm 28.98$ $43.80\pm 29.28$
Longbench ROUGE Score $16.70\pm 11.75$ $16.54\pm 11.82$ $17.50\pm 12.83$ $17.63\pm 12.81$
Appendix D Reproducibility and Resources
Inference was performed using the Hugging Face Transformers Wolf et al. (2020) and Accelerate Gugger et al. (2022) with FlashAttention Dao et al. (2022); Dao (2024). We integrated both Quanto and HQQ into the Language Model Evaluation Harness Gao et al. (2024) to enable systematic and reproducible evaluation of model performance under varying quantization schemes. Quantization error is measured by MSE, and confidence intervals use Bayesian variance. Hariri et al. (2025). Evaluations were executed on two High-performance Computing (HPC) clusters, detailed in Table 16.
Table 16: Specifications of Two High-Performance Computing (HPC) Clusters Used in This Study
Cluster Cluster A Cluster B
Processor AMD EPYC 7742 $\times$ 2 Intel Xeon Platinum 8468 $\times$ 2
RAM 2048 GB 2048 GB
GPU NVIDIA A100 $\times$ 8 NVIDIA H200 $\times$ 8
VRAM 80 GB HBM2e 141 GB HBM3e
Scale 5 nodes (40 A100) 5 nodes (40 H200)